OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

substitute sin(x) with u

OpenStudy (tyteen4a03):

Let sin(x) be y, solve the equation in terms of y then find the sine of the solutions.

OpenStudy (anonymous):

so the equation would be like this 7u^2 -14u +2 = -5

OpenStudy (anonymous):

u^2 -2u +1 =0

OpenStudy (anonymous):

solve for u

OpenStudy (anonymous):

then let sin(x) = u [you found (u) already ]

OpenStudy (anonymous):

solve for (x) then

OpenStudy (anonymous):

So, I got for u: (u-1)(u-1)... What do I do now... I'm confused. :-/ (and, thank you for the help!!)

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

so you have (u-1)^2 = 0

OpenStudy (anonymous):

that means u-1 = 0

OpenStudy (anonymous):

=> u = 1

OpenStudy (anonymous):

So, that is the solution for the equation?? I'm done?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

now Let Sin(x) = u = 1

OpenStudy (anonymous):

solve for (x)

OpenStudy (anonymous):

if Sin(x) = 1 then x= ......

OpenStudy (anonymous):

so x = 1/sin ??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

or simply x= pi/2

OpenStudy (anonymous):

x is an angle

OpenStudy (anonymous):

So, x = pi/2.... Are there any more solutions? I'm so sorry... I stink at math! :P

OpenStudy (anonymous):

i attached the complete solutions

OpenStudy (anonymous):

but simply u can say x = pi/2

OpenStudy (anonymous):

Oh, okay!! Thank you so much for your help!!! :)